On XYZ Island, there is an airport. The airport is the homebase of an unlimited number of identical airplanes. Each airplane has a fuel capacity to allow it to fly exactly 1/2 way around the world, along a great circle. The planes have the ability to refuel in flight without loss of speed or spillage of fuel. Though the fuel is unlimited, the island is the only source of fuel.What is the fewest number of aircraft necessary to get one plane all the way around the world assuming that all of the aircraft must return safely to the airport?

Grade:10

1 Answers

ADITI SINGH
7 years ago
Using the ideal gas law, a relationship between p and V is found to be p = nRT V . From the graph,it follows that the pressure is inversely proportional to the volume of the gas, e.g. p/p0 = 1/4 when V/V0 = 4. It can be seen from the equation above that this can only be true if the temperature of the gas remains constant. The curve OC is then representative of a isothermal process.